Joe needs 6 hours to file the day's paperwork. It take Charlotte 12 hours to do the same job. How long will it take if they work together?
Answer by saifoo.khan(162)

Hello,
1/6 + 1/12 = 1/x
Solve.
(2+1)/12 = 1/x
3/12 = 1/x
1/4 = 1/x
x = 4
So it will take both 4 hrs if they both are going to work.
THank you.


============================
If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($15/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($1.5-$3.0 per problem). Contact me at einstein.saifoo@gmail.com
Thank you. =)